if there is four girls and eight boys running for president or vice president of the student council what is the probability that two boys are elected and what is the probabilty that two girls are elected
Answer by stanbon(29625)

if there is four girls and eight boys running for president or vice president of the student council what is the probability that two boys are elected and what is the probabilty that two girls are elected
---------------------------
# of ways to pick 2 girls is 4C2 = [4*3]/[1*2] = 6
# of ways to pick 2 persons from the 10 is 10C2 = [10*9]/[1*2] = 45
------
Probability of electing 2 girls is 6/45 = 2/15
==================================================
Use the same counting procedure to get the following:
Probability of electing 2 boys is 8C2/10C2 = 28/45
===================================================
Cheers,
Stan H.